Before small batteries were available, carbide lamps were used for bicycle lights. Acetylene gas, C2H2, and solid calcium hydroxide were formed by the reaction of calcium carbide, CaC2, with water. The ignition of the acetylene gas provided the light. Currently, the same lamps are used by some cavers, and calcium carbide is used to produce acetylene for carbide cannons. What volume (in L) of C2H2 at 1.003 atm and 12.1°C is formed by the reaction of 15.57 g of CaC2 with water?
Answer #1

Molar mass of CaC2,

MM = 1*MM(Ca) + 2*MM(C)

= 1*40.08 + 2*12.01

= 64.1 g/mol

mass of CaC2 = 15.57 g

mol of CaC2 = (mass)/(molar mass)

= 15.57/64.1

= 0.2429 mol

Balanced chemical equation is:

CaC2 + 2 H2O -> C2H2 + Ca(OH)2

According to balanced equation

mol of C2H2 formed = moles of CaC2

= 0.2429 mol

Given:

P = 1.003 atm

n = 0.2429 mol

T = 12.1 oC

= (12.1+273) K

= 285.1 K

use:

P * V = n*R*T

1.003 atm * V = 0.2429 mol* 0.08206 atm.L/mol.K * 285.1 K

V = 5.67 L

Answer: 5.67 L